- 2x+4y+ 3Z = -27 xtsy-z = -2 3xty-42=25 Use Gauss-Bordan to solve write answer in format (o (xy, z 1
Answer #1

. - 24 +4 9 +3 2 - - 2 x +2 - Z = -2 3x +y42 스26 2 G 31-28 Augmented matvinis Ā = 51 -2.. 3 -리 25 Applying Gauss-jord an elim